Razor:

I did 20 + 1.50x / 59 If someone could check and quickly that'd be great

Razor:

1 attachment
dude:

I'm guessing you meant `=` instead of `/` [\(20 + 1.50x =59\)] In which case it is right (:

Razor:

Sweeet ty. Does that mean I can do 20 + 1.50, solve, then divide 59 by what's the answer of 20 + 1.50?

Razor:

Because that's the second part of the question

dude:

You can't combine terms that do not share the same variable \(20+1.50x=59\) (20 does not have x) So you would first be subtracting 20, then dividing by 1.50 [Because you want to get x by itself]

Razor:

59 - 20 = 39/1.50?

dude:

Yes that is the process I wouldn't write it like that because it would be confusing \(20+1.50x=59\) \(\color{red}{-20 }~~~~~~~~~~~~~~~\color{red}{-20}\) ==================== \(1.50x=39\) Then, \(1.50x=39\) \(\color{red}{\overline{~1.50~}}~~~~~\color{red}{\overline{1.50}}\) ==================== \(\large x=\frac{39}{1.50}\) [This can be simplified]

Razor:

Omg, thank you! Again, congrats on becoming a mod. Bye ;p

dude:

Ye that was right

Razor:

yeet oki <3